Point charges 3.5 μC and -2.4 μC are placed on the x axis at (11 m , 0) and (-11 m , 0), respectively

Find the point to the left of the negative charge where the electric potential vanishes.

X=
Answer #1

Let, the point is present at x to the left of the negative charge where the electric potential vanishes.

Electric potential is given by,

V = kq / r

so, kq1 / (d + x) - kq2 / x = 0

q1 / (d + x) = q2 / x

3.5*10^(-6) / (22 + x) = 2.4*10^(-6) / x

3.5x = 2.4 * (22 + x)

1.1x = 52.8

x = 48 m